OpenStudy (anonymous):

Find the roots of x^2+256=0 a.16+-i b.2+-4i c.+-16i d.+-4i

OpenStudy (atlas):

x^2 = -256 or, x^2 = 256 X -1 or, x = sqrt (256)X sqrt(-1)

OpenStudy (atlas):

I hope you know the sqrt of -1 is i and you can calculate the sqrt of 256

OpenStudy (anonymous):

so whats the answer?

OpenStudy (atlas):

what is the square root of 256??